OpenStudy (anonymous):

Factor the trinomial. w^3-5w^2-14w

OpenStudy (angela210793):

w^3-5w^2-14w=w(w^2-5w-14)

OpenStudy (anonymous):

\[w(w^2-5w-14)=w(w-7)(w+2)\]